OpenStudy (anonymous):

Factor the following expression into a product of two binomials: 25x^2 -90x +81

OpenStudy (anonymous):

25x^2 -90x +81 = (5x-9 )(5x-9)

OpenStudy (anonymous):

you are supposed to recognize this one as a "perfect square" \[(a-b)^2=a^2-2ab+b^2\] in this case \[a=5x,b=9\]

OpenStudy (anonymous):

Yes bradley I factored the other one for you. Just to let you know